kekeman:

A hot air balloon holds 4,328 cubic meters of helium. The density of helium is 0.1785 kilograms per cubic meter. How many kilograms of helium does the balloon contain, rounded to the nearest tenth of a kilogram? 3,747.9 kg 3,875.3 kg 772.5 kg 4,383.2 kg

AZ:

Density = mass/volume Density is 0.1785 kg/m^3 Volume is 4328 m^3 Mass is ?? We need to solve for mass

xxemilianaxx:

mass = density* volume

AZ:

0.1785 = x / 4328 if you multiply 4328 on both sides, what do you get? That's going to be the mass in kilograms

kekeman:

@az wrote:
0.1785 = x / 4328 if you multiply 4328 on both sides, what do you get? That's going to be the mass in kilograms
To get 9328

xxemilianaxx:

Or just do 4328 * 0.1785 = ? / since mass = density* volume

kekeman:

4328 * 0.1785

kekeman:

I got " 772.5 kg"

xxemilianaxx:

Perfect !

AZ:

That's correct!